# Enhancing Delta Compression in LLMs via SVD-based Quantization Error Minimization

### v2

@@ E.2 Inference Speed and Memory Cost
To demonstrate the impact of DeltaMix on inference speed and memory cost, we implement a simple Triton  (Tillet et al., 2019) kernel for DeltaMix. We compare our kernel with naive aligned models. Since there is no packing function of Delta-CoMe, we use our packing function and kernel for the Delta-CoMe method.
Following the setup in Yao et al. (2024), we assess the end-to-end system performance under varying arrival rates and request distributions. We consider two types of model popularity distribution: 1) Uniform: all models are equally popular. 2) Skewed: model popularity follows a Zipf-$\alpha$ distribution.
@@
DeltaMix 0.42 0.79 0.86 1.17
We evaluate the performance when serving 32 model variants of Qwen2.5-7B. Requests are sent to the serving system at a variable Poisson arrival rate ($\lambda$). To simplify, each request consists of 512 tokens, with the model generating one token as its response. We run the simulations for 100 seconds across different arrival rates and model distributions, measuring performance using two metrics: 1) end-to-end latency averaged over all requests; 2) Throughput, number of requests processed per second. All experiments are conducted on a single L40 GPU, with 28G of memory for storing models and the remaining memory for inference.
As shown in the Table 9, DeltaMix improves the throughput 6x and decreases end-to-end 100x compared to the naive method, because rather than loading the whole full-precision parameters, DeltaMix quantizes the delta-parameters so that a GPU can load more delta-parameters and switch them easily between CPU and GPU.
E.3 Time For Quantization
Simulation Optimization Quantization Total
Delta-CoMe Q_proj 0.0 0.0 3.6 50.5
K_proj 0.0 3.6
V_proj 0.0 3.6
O_proj 0.0 0.0 5.1
Up_proj 0.0 0.0 4.5
Gate_proj 0.0 4.5
Down_proj 0.0 0.0 25.6
DeltaMix Q_proj 4.7 8.5 0.5 143.6
K_proj 4.7 0.5
V_proj 4.7 0.5
O_proj 6.1 11.5 0.5
Up_proj 5.8 20.5 2.8
Gate_proj 5.8 2.8
Down_proj 30.2 22.5 11.0
In this section, we evaluate the quantization time of DeltaMix and Delta-CoMe within a single transformer block. The fundamental distinction between the two methods lies in their mixed-precision quantization strategies for each linear layer. DeltaMix determines the strategy by minimizing quantization loss, formulated as a 0/1 integer linear programming problem. To clarify the computational overhead, we decompose the quantization time into three components. The first is “simulation time”, which reflects the cost of estimating quantization loss under different bit-widths. The second is “optimization time”, incurred when solving the 0/1 integer linear programming problem. The third is the “quantization time” itself, representing the cost of quantizing each linear layer according to the selected strategy. The corresponding results for one transformer block of Qwen2.5-Math-7B-Instruct, which contains 28 blocks in total, are summarized in Table 10. For Delta-CoMe, both simulation and optimization times are zero because its mixed-precision quantization strategy is predetermined and applied uniformly across all linear layers; consequently, the entire forward pass is accounted for within the quantization time. In contrast, DeltaMix incurs additional simulation and optimization costs, which are higher for Up_proj, Gate_proj, and Down_proj due to their larger row or column dimensions. Specifically, simulation time increases with the number of columns, while optimization time grows with the number of rows. Notably, DeltaMix ’s quantization time is shorter than that of Delta-CoMe, since the forward pass is already included in its simulation stage.
Overall, although DeltaMix takes 3x more time than Delta-CoMe, it only requires 1.2 hours for 7B models and 2.4 hours for 14B models on a single L20 GPU, which is acceptable. In contrast to Delta-CoMe’s degraded performance on the large norm of $\Delta\mathbf{W}$, DeltaMix consistently achieves comparable or better results across all scenarios.
E.4 Analyzing the Bit Allocation Results
@@
Delta-CoMe (Ping et al., 2024) empirically posits that singular vectors corresponding to larger singular values are more significant and, therefore, necessitate higher-bit representations. We further examine whether DeltaMix adheres to this assumption, specifically by using singular values alone to evaluate importance. We compute the Kendall rank correlation coefficient $\tau$, between the bit sequence and the singular value sequence for each $\mathbf{W}$. The coefficient is a measure of rank correlation, ranging from -1 to 1, reflecting the similarity of the orderings of the data when ranked by each of the quantities. If the method strictly adhered to the assumption of using singular values alone for importance assessment, singular vectors with larger singular values would always receive higher bit-width, resulting in a consistent $\tau=1$ across all $\mathbf{W}$. However, for the DeepSeek-R1-Distill-Qwen-7B model with DeltaMix, we observe a $\tau$ of 0.95 for the $\mathbf{W}$ of the key projection at layer 28. This indicates that DeltaMix goes beyond singular values, taking into account both the “scaling” term and the “difference” term.
E.5 Analyzing the Quantization Error Across Weight Types and Layers
@@
DeltaMix 1.83 1.74 1.36 1.59 6.58 8.89 DeltaMix 2.28 2.22 1.57 0.59 6.65 2.07
Param Down_proj Param Average

### v3

@@ C.6 Inference Speed and Memory Cost
To demonstrate the impact of PrinMix on inference speed and memory cost, we implement a simple Triton  (Tillet et al., 2019) kernel for PrinMix. We compare our kernel with naive aligned models. Since there is no packing function of Delta-CoMe, we use our packing function and kernel for the Delta-CoMe method.
Following the setup in Yao et al. (2024), we assess the end-to-end system performance under varying arrival rates and request distributions. We consider two types of model popularity distribution: 1) Uniform: all models are equally popular. 2) Skewed: model popularity follows a Zipf-$\alpha$ distribution. We evaluate the performance when serving 32 model variants of Qwen2.5-7B. Requests are sent to the serving system at a variable Poisson arrival rate ($\lambda$). To simplify, each request consists of 512 tokens, with the model generating one token as its response. We run the simulations for 100 seconds across different arrival rates and model distributions, measuring performance using two metrics: 1) end-to-end latency averaged over all requests; 2) Throughput, number of requests processed per second. All experiments are conducted on a single L40 GPU, with 28G of memory for storing models and the remaining memory for inference.
As shown in the Table 14, PrinMix improves the throughput 6x and decreases end-to-end 100x compared to the naive method, because rather than loading the whole full-precision parameters, PrinMix quantizes the delta-parameters so that a GPU can load more delta-parameters and switch them easily between CPU and GPU.
@@
PrinMix 0.42 0.79 0.86 1.17
C.7 Analyzing the Bit Allocation Results
@@
Delta-CoMe (Ping et al., 2024) empirically posits that singular vectors corresponding to larger singular values are more significant and, therefore, necessitate higher-bit representations. We further examine whether PrinMix adheres to this assumption, specifically by using singular values alone to evaluate importance. We compute the Kendall rank correlation coefficient $\tau$, between the bit sequence and the singular value sequence for each $\mathbf{W}$. The coefficient is a measure of rank correlation, ranging from -1 to 1, reflecting the similarity of the orderings of the data when ranked by each of the quantities. If the method strictly adhered to the assumption of using singular values alone for importance assessment, singular vectors with larger singular values would always receive higher bit-width, resulting in a consistent $\tau=1$ across all $\mathbf{W}$. However, for the DeepSeek-R1-Distill-Qwen-7B model with PrinMix, we observe a $\tau$ of 0.95 for the $\mathbf{W}$ of the key projection at layer 28. This indicates that PrinMix goes beyond singular values, taking into account both the “scaling” term and the “difference” term.
C.8 Analyzing the Quantization Error Across Weight Types and Layers
@@
PrinMix 1.83 1.74 1.36 1.59 6.58 8.89 PrinMix 2.28 2.22 1.57 0.59 6.65 2.07
Param Down_proj Param AVG